(x+y)2+(x-3y)2=: (x+y)2+(x-3y)2=
Answer by orca(336)

First use the formula
(a+b)^2=a^2+2ab+b^2
and
(a-b)^2=a^2-2ab+b^2
to expand,
then combine like terms.
(x+y)^2+(x-3y)^2
= x^2+2xy+y^2+x^2-2x*(3y)+(3y)^2
= x^2+2xy+y^2+x^2-6xy+9y^2
= 2x^2-4xy+10y^2